# ACORD: An Expert-Annotated Retrieval Dataset for Legal Contract Drafting

### v1

@@ 5 Discussion and Future Work
Lexical Matching Difficulties
Some queries perform better with the MiniLM bi-encoder retriever than with BM25. For example, BM25 scores 0 on all metrics for the “as-is clause” query (see Table E-2) and near-zero scores even when paired with GPT-4o as the reranker. The reranking does not help much because BM25 does not retrieve relevant clauses in the top 100 results; the reranker only operates on the top 100 results. Hence, the recall@100 would be low or even zero. The dense retriever MiniLM bi-encoder performs much better than BM25 on the “as-is clause" query and many other queries.
@@ 6 Conclusion
We introduced ACORD, the first information retrieval dataset for contract drafting fully annotated by experts. ACORD aims to facilitate research on clause retrieval, a foundational contract drafting task focusing on the most complex and heavily negotiated clauses in commercial contracts. ACORD includes 114 queries and over 126,000 query-clause pairs, each ranked on a scale from 1 to 5 stars. We evaluated 20 retrieval methods on ACORD and found that performance is promising but still has significant room for improvement. We also showed that model design markedly influences performance, suggesting that algorithmic improvements from the NLP community will help solve this challenge.
@@ A.2 Annotator Recruitment and Demographics
The annotators were recruited by a nonprofit organization for the purpose of creating ACORD. The lawyer annotators are volunteers and do not receive compensation. Student annotators receive hourly compensation. All annotators were informed of the purpose of their involvement and consented to the inclusion of their annotations in ACORD.
@@ C.2 Ethics Board Review
This research project was reviewed for ethics considerations by the the ETH Zürich Ethics Commission.
@@ C.4 Clause Extraction Methodology in Clause Corpus
Figure C-1: Histogram over the number of words in 2- to 5-star clause in the ACORD dataset. The figure also includes statistics showing that 2- to 5-star clauses have an average of 225 words and a median of 146 words in each clause.
@@ F.1 Performance Metrics
Performance is measured using standard metrics like NDCG@5, NDCG@10, 5-star-precision@5, 4-star-precision@5, and 3-star-precision@5. The NDCG@k measures the normalized discounted cumulative gain of the top k returned results. We use the standard definition of NDCG@k seen in (1).

### v2

@@ 6 Conclusion
We introduced ACORD, the first expert-annotated clause retrieval dataset for contract drafting tasks. ACORD aims to facilitate research on clause retrieval, a foundational contract drafting task focusing on the most complex and heavily negotiated clauses in commercial contracts. ACORD includes 114 queries and over 126,000 query-clause pairs, each ranked on a scale from 1 to 5 stars. We evaluated 20 retrieval methods on ACORD and found that performance is promising but still has significant room for improvement. We also showed that model design markedly influences performance, suggesting that algorithmic improvements from the NLP community will help solve this challenge.
@@ A.2 Annotator Recruitment and Demographics
A nonprofit organization recruited the annotators to create ACORD. The lawyer annotators are volunteers and do not receive compensation. Student annotators receive hourly compensation. All annotators were informed of the purpose of their involvement and consented to the inclusion of their annotations in ACORD.
@@
A.3 Annotator Roles and Contributions
Our annotation team consisted of 12 experienced attorneys and 10 students working under the attorneys’ direct supervision following 5-10 hours of training. As detailed in Section 3 (“Labeling Process”), each clause pair in the dataset was independently annotated by three individuals: one student annotator and two experienced attorneys. In instances of disagreement, annotations were reviewed and adjudicated by a panel comprising three to six experienced lawyers to ensure consistency and legal accuracy. Annotator disagreement was tracked by the number of scores revised during this reconciliation step.
A substantial proportion of 1-star clauses in the dataset originated from data augmentation using the CUAD dataset.
All attorney contributions were provided on a pro bono basis. Student annotators received hourly compensation ranging from \$20 to \$35.
@@ C.2 Ethics Board Review
This research project was reviewed for ethics considerations by the REDACTED FOR ANONYMITY.
@@ C.4 Clause Extraction Methodology in Clause Corpus
Figure C-1: Histogram over the number of words in 2- to 5-star clauses in the ACORD dataset. The figure also includes statistics showing that 2- to 5-star clauses have an average of 225 words and a median of 146 words in each clause.
@@ F.1 Performance Metrics
Performance is measured using standard metrics like NDCG@5, NDCG@10, 5-star-precision@5, 4-star-precision@5, and 3-star-precision@5. When computing the metrics, the scores are labeled from 0 to 4 rather than 1 to 5 to ensure irrelevant results (1-star clauses) are weighted appropriately.
The NDCG@k measures the normalized discounted cumulative gain of the top k returned results. We use the standard definition of NDCG@k seen in (1).